OpenStudy (anonymous):

Solve on the interval 0,2pi) 1+sin theta=1/2

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

that means \(\sin(\theta)=-\frac{1}{2}\)

OpenStudy (misty1212):

find the numbers in one rotation of the unit circle that have \(-\frac{1}{2}\)as the second coordinate there are two of them right?

OpenStudy (anonymous):

uhh 7pi/6 and 11pi/6? @misty1212